OpenStudy (anonymous):

Which expression is equivalent to (82)−4?

OpenStudy (anonymous):

is this supposed to be a multiple choice question?

OpenStudy (astrophysics):

That's just 82-4..

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how can I show the answers?

OpenStudy (astrophysics):

We don't need them, 

OpenStudy (astrophysics):

Do you know how to subtract?

OpenStudy (anonymous):

(8^2) -4 is the question

OpenStudy (astrophysics):

Oooh!

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

what's 8 squared

OpenStudy (astrophysics):

\[(8^2)-4 \implies 8 \times 8 - 4\]

OpenStudy (astrophysics):

\[8^2 = 8 \times 8 \]

OpenStudy (anonymous):

so it would be 8^2 ?

OpenStudy (astrophysics):

What?

OpenStudy (astrophysics):

I'm saying 8^2 is 8 times 8

OpenStudy (astrophysics):

x^2 = x * x

OpenStudy (anonymous):

its a multiple choice question with A,b, c, and d

OpenStudy (astrophysics):

I don't care about the multiple choice

OpenStudy (anonymous):

do you know what 8 times 8 is??

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Just answer the question what is \[8 \times 8 -4\]

OpenStudy (anonymous):

60

OpenStudy (astrophysics):

Yeah, that's it.

OpenStudy (anonymous):

thank you

OpenStudy (astrophysics):

Just remember this, if you see something squared, cubed, what ever, it just means that it's being multiplied by itself by the amount of the exponent.

OpenStudy (anonymous):

hmm.. thank you for your help (:

OpenStudy (astrophysics):

So \[8^5 \implies 8 \times 8 \times 8 \times 8 \times 8\]

OpenStudy (astrophysics):

Np :P

OpenStudy (anonymous):

(82)−4? 82 x -4